OpenStudy (arindameducationusc):

If x is a positive number such that 2x+12 is perfectly divisible by x, then the number of possible values of x is

myininaya (myininaya):

so you are given that (2x+12)/x is an integer

myininaya (myininaya):

my first idea is let (2x+12)/x=k where k is that integer

myininaya (myininaya):

play with this equation

myininaya (myininaya):

big hint: group all your x terms on one side

myininaya (myininaya):

oops plus get rid of the division part :p

OpenStudy (arindameducationusc):

I got it

OpenStudy (arindameducationusc):

answer is 6

OpenStudy (arindameducationusc):

another question

myininaya (myininaya):

that's right

OpenStudy (arindameducationusc):

Let K be a positive integer such that k+4 is divisible by7. Then the smallest positive integer n, greater than2, such that k+2n is divisible by 7 equals?

ganeshie8 (ganeshie8):

k + 4 is divisibly by 7 implies there exists an integer 'a' such that k + 4 = 7*a

OpenStudy (arindameducationusc):

ok

ganeshie8 (ganeshie8):

k + 2n is divisibly by 7 implies there exists an integer 'b' such that k + 2n = 7*b

ganeshie8 (ganeshie8):

solve 'k' from first equation and plug it in the second equation

OpenStudy (arindameducationusc):

oh

OpenStudy (arindameducationusc):

the 2 is given for creating a confusion?

ganeshie8 (ganeshie8):

We're not done yet

OpenStudy (arindameducationusc):

oh

ganeshie8 (ganeshie8):

Have you plugged the expression for 'k' in the second equation ?

OpenStudy (arindameducationusc):

i get a=b

OpenStudy (arindameducationusc):

7a-4+4=7b

ganeshie8 (ganeshie8):

How has 2n turned into 4 ?

OpenStudy (arindameducationusc):

oops,wait

OpenStudy (arindameducationusc):

7a-4+2n=7b

ganeshie8 (ganeshie8):

You have so far : k + 4 = 7*a k + 2n = 7*b solving 'k' from first equation and plugging it in second equation gives 7a - 4 + 2n = 7b

ganeshie8 (ganeshie8):

Yes, rearrange a bit and get -4 + 2n = 7(b-a)

ganeshie8 (ganeshie8):

Since right hand side is divisible by 7, we must choose "n" such that the left hand side is also divisible by 7

ganeshie8 (ganeshie8):

For what values of "n" is the expression "-4 + 2n" divisible by 7 ?

OpenStudy (arindameducationusc):

hmmm

OpenStudy (arindameducationusc):

thinkin, wait

ganeshie8 (ganeshie8):

Plugin values like n = 0,1,2,3, etc and see which one makes the expression -4 + 2n divisible by 7

OpenStudy (arindameducationusc):

9

ganeshie8 (ganeshie8):

good, how about n=2

OpenStudy (arindameducationusc):

it becomes zero

ganeshie8 (ganeshie8):

Is 0 divisible by 7 ?

OpenStudy (arindameducationusc):

7*0=0, so ya

ganeshie8 (ganeshie8):

Nice, notice that 9 = 2+7

ganeshie8 (ganeshie8):

Can you guess another value for n that makes the expression -4 + 2n divisible by 7 ?

OpenStudy (arindameducationusc):

yes

ganeshie8 (ganeshie8):

do it

OpenStudy (arindameducationusc):

16

ganeshie8 (ganeshie8):

how ?

OpenStudy (arindameducationusc):

i will tell you,

OpenStudy (arindameducationusc):

-4+2n=28

OpenStudy (arindameducationusc):

n=16

OpenStudy (arindameducationusc):

21 is not possible, so i took 28

ganeshie8 (ganeshie8):

Notice again 16 = 2 + 7*2

ganeshie8 (ganeshie8):

See the pattern ?

ganeshie8 (ganeshie8):

Any integer of form " 2 + 7*t " makes the expression "-4 + 2n" divisible by 7

ganeshie8 (ganeshie8):

For example, n = 2 + 7*3 works just fine

OpenStudy (arindameducationusc):

ok, cool

OpenStudy (arindameducationusc):

thanks. Your motivation encourages me to solve more problems. If I get problems I will let you know :)

ganeshie8 (ganeshie8):

Np :) See if you can figure out why any integre of form "2 + 7*t" makes the expression "-4+2n" divisible by 7

ganeshie8 (ganeshie8):

when you're free

OpenStudy (arindameducationusc):

honestly I did by counting and checking if it was divisible by 7

OpenStudy (arindameducationusc):

putting the values 0,1,2,3,4....

OpenStudy (arindameducationusc):

that pattern.. man, you are a genius. How did you take out that!

OpenStudy (arindameducationusc):

I think you just put it in format 2+7t after placing the n=0,1,2,3,4, so on

OpenStudy (robtobey2):

3,4,5,6,8,14 Refer to the Mathematica attachment.